# Challenges and Opportunities for Visual Analytics in Jurisprudence

### v2

@@ 5.1.1 $W_{\texttt{P1}}$: Discovery and Scoping
Next, Alice evaluates the documents in the search results ($E_{2}$). The results are sorted by relevance, but the platform does not make the ranking criteria transparent, leaving Alice uncertain about the prioritization (Inefficiency of Current Interfaces). She skims through the initial results and discovers that some relevant documents are inaccessible due to the limits of her subscription (Accessibility Issues). Despite their potentially high relevance, Alice is forced to exclude these documents from her analysis for now. The first results predominantly include journal articles, court rulings, and commentaries rather than primary legal statutes. To supplement this, Alice manually expands her search to explanatory memoranda in other databases ($E_{4}$), such as the Dokumentations- und Informationssystem für Parlamentsmaterialien (DIP) [94] (Navigational Complexity). Recognizing the importance of these documents for contextual analysis, she notes them for detailed inspection in $W_{\texttt{P2}}$.
5.1.2 $W_{\texttt{P2}}$: Analysis and Interpretation
@@
In the second phase ($W_{\texttt{P2}}$), Alice analyzes the documents identified in the first phase $W_{\texttt{P1}}$ ($T_{2}$). She managed to get some documents she could not access online as prints from her university library. Alice then scanned and saved them as digital documents. She focuses on connecting the legal norms highlighted by beck-online – such as Art. 3 GG, § 23 BPolG, and Art. 14 EMRK – with the content of the selected documents ($E_{3}$). beck-online allows Alice to preview these norms but does not support visualizing relationships between them. Alice draws on her expertise to recall that § 23 BPolG, which permits identity checks by the federal police, is conceptually linked to §§ 2-4 BPolG, which outline the responsibilities of the federal police (Reliance on Tacit Knowledge). However, beck-online lacks tools to map or visualize such connections explicitly, requiring Alice to track the relationships as she progresses through her analysis mentally (Navigational Complexity).
@@
She begins by examining the first matching document from a journal. Due to its limited structural organization, Alice skims the text to identify relevant legal norms and notes them mentally ($E_{2}$). This process repeats for the remaining documents, including those accessed through supplemental searches in other databases. Alice consults their details whenever encountering unfamiliar norms by following backlinks in the legal database. Through her analysis, Alice identifies arguments supporting her assumption that German law inadequately prevents racial profiling. However, eliciting these insights requires substantial manual effort as she digs through multiple documents, particularly when integrating the scanned paper documents. Alice also suspects that norms related to the state police laws (such as the Polizeiaufgabengesetz (Police Tasks Act) in Bavaria) might offer additional insights. However, she finds no references to these in her current search results, indicating a gap in the coverage (Accessibility Issues).
5.1.3 $W_{\texttt{P3}}$: Synthesis and Documentation
[…]
@@ 6 Visual Analytics for Jurisprudence
One can categorize the points raised by the interviewees into several distinct and recurring challenges faced by legal scholars when analyzing and applying legal materials:
@@
    Accessibility Issues: Licensing restrictions and lack of digital availability limit access to essential legal commentaries and court rulings, particularly at lower court levels. Participants emphasized the importance of data accessibility and availability of digitized legal texts in streamlining their work. For instance, $E_{8}$ noted that access to digital resources accelerates research ($T_{1}$), whereas manual retrieval of printed materials causes significant delays ($E_{5}$). However, the participants pointed out that the duopoly of beck-online and juris restricts access to crucial documents. Further, lower court rulings, in particular, remain difficult to obtain due to data protection regulations [25].
@@
    Navigational Complexity: A lack of approaches for navigating relationships between legal texts to identify relevant norms. Participants often rely on manual methods to understand and navigate the formal legal relationships between documents ($T_{2}$). These methods include analyzing tables of contents and formal document structures ($E_{7}$ and $E_{9}$), exploring explanatory memoranda for context ($E_{4}$ and $E_{7}$) and using hyperlink-based backlinks between related documents, or manually searching for connections when links were unavailable ($E_{2}$). Those workflows demand significant manual effort, underscoring the need for tools that enable exploratory navigation of document hierarchies and relationships.
@@
[…]
@@
    AI Usage: A cautious but growing interest in AI-based tools is contingent on their reliability and explainability. However, opinions on AI and LLMs vary: while $E_{6}$ rarely uses LLMs, $E_{2}$ and $E_{7}$ highlight concerns about transparency and reliability. Some participants ($E_{2}$) use LLMs to filter search results retrieved from commercial platforms ($T_{1}$), suggesting cautious optimism about the potential of LLMs.
@@
    Reliance on Tacit Knowledge: The relationships between legal materials are essential for sophisticated legal reasoning. However, many of these relationships are not explicitly encoded in legal documents. Instead, scholars must draw on their domain expertise, which includes implicit relationships to interpret norms. Integrating their implicit knowledge with the explicit knowledge conveyed in current interfaces is labor-intensive and error-prone.
Apart from the licensing issues, these findings emphasize a need for more adequate tooling support in the legal domain. VA offers ways to address these challenges through collaboration between humans and AI by enabling semantic exploration and systematic analysis of legal texts. The individual phases in the workflow (refer to subsection 5.1) need different types of support for these challenges. As discussed previously (refer to section 2), some of the challenges align with standard practices in text-based research domains, which allows us to reuse some existing concepts, but with the need to adapt and integrate them to the specific needs of law scholars.
In computer science, a separation of concerns is fundamental to the architecture of our modern computers. It separates input and output from data management and representation, and algorithmic instructions [95]. In recent interpretation, VA also considers the interaction of humans with computers from three perspectives: data, model, and visualization [47]. Data describes the information related to the analytical problem, which, in law, is not only the raw text but also their interdependency and metadata. A model can refer to descriptive statistics or a sophisticated computational algorithm that processes the data, which is akin to the cognitive capabilities of human legal reasoning. Visualization refers to the visual representation of data and models, which can be a line-by-line block of text on a display in a specific formatting or a flow chart of causal relationships between norms in the legal domain. These aspects frame the three related but distinct perspectives we apply to the recurring challenges previously identified as leverage points [93].
We consider the following aspects to contextualize the three perspectives with the identified challenges and the legal domain. User interfaces of legal databases visualize document collections and enable their navigation, but are currently complex and inefficient. The complexity results from intricate formal legal structures, and the inefficiencies are due to the lack of contextual and semantic understanding in syntactic keyword searches. VA can support this navigation of the information space (refer to subsection 6.1). The explicit and implicit knowledge encoded in legal documents through references and semantic formulations must be sufficiently represented (refer to subsection 6.2) in the form of data structures and relations. Since tacit knowledge is not readily encoded in the data, it is also partly responsible for the current inefficiency of user interfaces and the difficulties in modeling this information. The logical analysis and reasoning, a manual process so far, can be supported through the navigation and tailored interfaces by making relevant information more accessible and connected (refer to subsection 6.3). In all three perspectives, AI solutions can be employed to support data filtering, to improve the context-dependent prediction of relevant documents, and to help with analytical reasoning, for example.
@@
Based on the computer science background and the mapping from challenges to core topics, we identify three key areas of VA that hold significant benefits for jurisprudence: (1) data navigation: for effective exploration and cross-connection of document collections, (2) knowledge representation: for properly structuring data and metadata while integrating domain knowledge with explicit knowledge from legal entities, and (3) analytical reasoning: for synthesizing and documenting findings with the support of AI. In Figure 2, we showcase the visual components of an exemplary traditional user interface of a commonly used German legal database. Alongside this traditional UI, we propose potential enhancements through VA techniques. The traditional keyword-based search (refer to section 4) consists of three parts: (1) the formal legal structure of the inspected law for data navigation as a hierarchical list, (2) the display of the legal norm’s contents as a main body of text, and (3) a simple list of related legal documents. Based on the literature and our findings, the following enhancements are possible. On the left-hand side, additionally to the hierarchical list, the UI (1) visualizes the hierarchy of the Bundespolizeigesetz (Federal Police Act) using a collapsible tree structure as a conceptual tree map. It contains the law’s formal structure labeled by its section titles. Users can navigate the law through interaction with the structure, moving between legal norms. The UI (2) renders the selected legal norm’s contents at the center, representing explicit knowledge. The rendering highlights the keywords from the query in the contents verbatim. The user interface backlinks to their documents for explicitly referenced legal entities, simultaneously displaying a preview on hover. On the right-hand side, the legal database (3) lists legal documents related to the currently displayed document, like legal norms, law commentary, and court rulings. For the three key areas of VA discussed above, we analyze the deficiencies, explore the state-of-the-art literature, and distill possible improvements in jurisprudence through VA while discussing the intersection’s challenges.
@@ 6.1 Data Navigation
In VA, data navigation encompasses the exploration of and the movement in information spaces [58]. For jurisprudence, this translates to navigating formally structured legal document collections. The traditional UI of legal databases (refer to Figure 2) relies on a collapsible tree-based list to represent hierarchies. However, this approach fails to provide semantic context during navigation. Since legal structures can be nested several levels deep [49, p. 152], their tree representation can become cumbersome. Additionally, tree-based lists inadequately support the simultaneous navigation of multiple legal structures – a common requirement in jurisprudential research.
@@
Merkl and Schweighofer introduced the concept of navigating such corpora using a Hierarchical Feature Map [68], a Treemap variant that organizes clusters across layers representing the hierarchy. Each layer corresponds to a hierarchical level, with subsequent layers revealing sub-clusters. Despite their improvement, the static, semi-automatically generated maps require manual input for cluster titles, limiting scalability. Building on this idea, Lettieri et al. started employing Treemaps [96] in the Legal Doctrine Semantic Navigator for navigating formally structured legal documents through drill-down and roll-up operations [74]. Later, they introduced the sliding Treemap, a variation, enabling the navigation of rulings by the European Court of Justice and other legal documents on mobile devices [97]. Treemaps with interactive drill-downs can display multiple paths simultaneously across hierarchies and can visually encode guidance by enlarging documents of higher relevance [74] (refer to Figure 2, a).
@@
Beyond Treemaps, Icicle Plots [98] have emerged as an effective visualization method for hierarchical data. DH researchers employ these plots for document comparison [99]. Kim et al. use Icicle Plots to visualize a hierarchy of topics during their interactive modeling [100], while Tytarenko et al. also visualize a hierarchy of topics to compare documents [99] (refer to Figure 2, a).
@@
We propose a hybrid approach combining the strengths of both visualization techniques based on the following observations. Since Treemaps perform worse at displaying large hierarchies than Icicle Plots [101], distributing the visual representation between these complementary visualizations is beneficial. A Treemap would display the first few layers of the currently selected path as an overview, while a juxtaposed Icicle Plot would present the remainder of the path in the hierarchy. Employing the two visualizations in parallel mitigates scalability issues and leverages the benefits of interactively linking both visualizations [102], offering complementary perspectives on legal corpora. This combination would support legal scholars during the initial workflow phase (refer to $W_{\texttt{P1}}$), facilitating the discovery and scoping of legal document collections.
@@
Alternative hierarchical visualization techniques such as Bubble Treemaps or Sunbursts [101] can also be employed but perform worse for small hierarchies than our proposed approaches. Still, we emphasize the need for exploring other design alternatives such as circular grid layouts [103] and hierarchical edge bundling [104] for radial visualizations like a radial dendrogram. Though less prevalent in the legal domain, these alternatives are worth considering for expanding the design space.
@@ 6.2 Knowledge Representation
The interactions with visualizations for data navigation heavily depend on the user’s prior knowledge, explicit and tacit [48]. In jurisprudence, explicit knowledge can manifest as connections between legal documents [75], where one norm references another legal norm, establishing a relationship. The traditional UI of legal databases displays these relationships through backlinks, with a popup and metadata, thereby representing explicit knowledge. The tree-based list of the law’s hierarchy to the left and the list of related legal documents on the right also express explicit knowledge contained in legal documents. Meanwhile, legal scholars’ tacit knowledge about the applications of these legal norms is internalized as part of legal scholarship. However, it is not articulated [4], which would be especially important for legal reasoning (refer to subsubsection 5.1.2).
@@
In the literature, various types of visualizations, such as stacked graphs and geographic maps, have been employed to visualize the metadata of legal documents [70]. Similarly, Resck et al. visualize the structure of legal documents by semantically segmenting them into paragraphs [76]. Meanwhile, other works statically analyze the complexity of legal documents by visualizing explicit references between their articles [72]. Beyond jurisprudence, explicit knowledge of text-based documents is visualized in the context of discussion forum posts [105] and document comparison [99]. However, these examples, both within and outside jurisprudence, often neglect to facilitate the articulation of tacit knowledge. Instead, VA should consider tacit knowledge [45]. For instance, Wagner et al. allow experts for cybersecurity to classify behavioral rules of software for identifying malware [61], while Mistelbauer et al. enhance medical visualizations by ranking views based on the experience of medical experts [106]. Endert et al. define several semantic interactions that support analytical reasoning for text analytics [107]. The authors also provide an implementation, letting users determine the importance of phrases through text highlighting. Accordingly, documents containing the exact phrase become closer in a spatial layout (refer to Figure 2, b). Similarly, a diagram of concepts connected by labeled arcs constitutes a Concept Map, where the arcs created by semantic interactions describe tacit knowledge [108] (refer to Figure 2, b).
@@
Translating these solutions from VA to jurisprudence, we observe that visually laying out legal concepts in a map is an established approach [92][109, p. 48][110]. Concept maps can enhance VA for jurisprudence by converting legal scholars’ domain knowledge into explicit knowledge through semantic interactions, thereby improving the effectiveness of LLMs. With KMTLabeler, Wang et al. propose a labeling tool for medical texts integrating VA with LLMs [111]. This tool allows experts to express their domain knowledge by adjusting keyword weights and defining labeling rules. The resulting explicit knowledge is incorporated into the embedding generation of the LLM and used to improve the labeling of the medical texts. We can also connect the knowledge representation to legal reasoning since a Concept Map resembles a network. Both support the latter two phases of the jurisprudence workflow (refer to $W_{\texttt{P2}}$ and $W_{\texttt{P3}}$).
@@
While networks are suitable for supporting this workflow phase, node-link diagrams may not be the most effective visualization method. Matrices, i.e., tabular displays of graphs, are proven methods for visualizing network data [112]. However, they are underutilized in VLA. Hence, studying hybrid approaches and alternative representations specific to the legal domain can enrich the VLA design space [112].

### v3

@@ 5.1 $W_{\texttt{P1}}$: Discovery & Scoping
Next, Alice evaluates the documents in the search results ($E_{2}$). The results are sorted by relevance, but the platform does not make the ranking criteria transparent, leaving Alice uncertain about the prioritization (Inefficiency of Current Interfaces). She skims through the initial results and discovers that some relevant documents are inaccessible due to the limits of her subscription (Accessibility Issues). The first results predominantly include journal articles, court rulings, and commentaries rather than primary legal statutes. To supplement this, Alice manually expands her search to explanatory memoranda in other databases ($E_{4}$), such as the Dokumentations- und Informationssystem für Parlamentsmaterialien (DIP) [deutscherbundestagDIPDokumentationsUnd] (Navigational Complexity). Recognizing the importance of these documents for contextual analysis, she notes them for detailed inspection in $W_{\texttt{P2}}$.
5.2 $W_{\texttt{P2}}$: Analysis & Interpretation
@@
In the second phase ($W_{\texttt{P2}}$), Alice analyzes the documents identified in the first phase $W_{\texttt{P1}}$ ($T_{2}$). She focuses on connecting the legal norms highlighted by beck-online – such as Art. 3 GG, § 23 BPolG, and Art. 14 EMRK – with the contents of the selected documents ($E_{3}$). beck-online allows Alice to preview these norms, but does not support visualizing relationships between them. Alice draws on her expertise to recall that § 23 BPolG, which permits identity checks by the federal police, is conceptually linked to §§ 2-4 BPolG, which outline the responsibilities of the federal police (Reliance on Tacit Knowledge). However, beck-online lacks tools to map or visualize such connections explicitly, requiring Alice to track the relationships as she progresses through her analysis mentally (Navigational Complexity).
@@
She begins by examining the first matching document from a journal. Due to its limited structural organization, Alice skims the text to identify relevant legal norms and notes them mentally ($E_{2}$). This process repeats for the remaining documents, including those accessed through supplemental searches in other databases. Alice consults their details whenever encountering unfamiliar norms by following backlinks in the legal database. Through her analysis, Alice identifies arguments supporting her assumption that German law inadequately prevents racial profiling. However, eliciting these insights requires substantial manual effort as she digs through multiple documents. Alice also suspects that norms related to the state police laws (such as the Polizeiaufgabengesetz (Police Tasks Act) in Bavaria) might offer additional insights. However, she finds no references to these in her current search results, indicating a gap in the coverage (Accessibility Issues).
5.3 $W_{\texttt{P3}}$: Synthesis & Documentation
In the final phase ($W_{\texttt{P3}}$), Alice synthesizes her findings and documents the outcomes of her analysis ($E_{4}$$T_{3}$). She integrates her mental map of relationships between legal norms with the arguments and insights derived from the reviewed documents. She organizes these findings into a coherent narrative using a word processing program, ensuring that references to relevant legal texts and scholarly arguments support her conclusions. Alice imports the key documents into her reference management software for citation and further reference ($E_{2}$). This phase remains labor-intensive due to the absence of automated insights from multiple sources that visualize the relationships between legal norms (Navigational Complexity). Despite these limitations, Alice completes the synthesis and prepares to revisit it in subsequent iterations, aiming to address gaps identified in previous phases.
@@ 6 Visual Analytics for Jurisprudence
We group the interviewees’ comments into recurring challenges faced by legal scholars when analyzing and applying legal materials:
@@
    Accessibility Issues: Limited digital access and licensing restrictions hinder the use of commentaries and rulings, especially from lower courts. Participants ($E_{8}$) stressed that data availability accelerates research ($T_{1}$), while relying on print sources causes delays ($E_{5}$). They also noted that the duopoly of beck-online and juris restricts access, and that lower-court rulings are often inaccessible due to data protection rules [ltoTransparenzJustizStagnation].
@@
    Navigational Complexity: Navigating relationships between legal texts requires substantial manual effort. Participants reported using tables of contents ($E_{7}$, $E_{9}$), explanatory memoranda ($E_{4}$, $E_{7}$), or manual searches when links were missing ($E_{2}$). They emphasized the need for tools that better support exploratory navigation of document hierarchies and semantic relationships ($T_{2}$).
@@
    Inefficiency of Current Interfaces: Existing platforms (e.g., beck-online, juris) emphasize syntactic rather than semantic search. Participants cited limited support for domain-specific language ($E_{4}$, $E_{6}$, $E_{7}$), failure to find implicit legal concepts ($E_{3}$), and over-reliance on keyword matching ($E_{4}$$T_{3}$). To cope, some ($E_{2}$, $E_{4}$, $E_{5}$) used browser search within results, underscoring the demand for features such as fuzzy or semantic search and contextualization ($T_{S}$, $E_{6}$).
@@
    AI Usage: Interviewees showed cautious interest in AI tools, stressing the need for reliability and transparency. While $E_{6}$ rarely used LLMs, $E_{2}$ and $E_{7}$ raised concerns about their trustworthiness. $E_{2}$ reported using LLMs to filter search results ($T_{1}$), suggesting guarded optimism about their potential.
@@
    Reliance on Tacit Knowledge: Legal reasoning depends on relationships between documents, but many links remain implicit. Scholars rely on their domain expertise to infer connections, a process they described as labor-intensive and error-prone. They identified value in tools that could better integrate explicit and tacit knowledge, which is currently labor-intensive and error-prone.
Apart from open data, these findings emphasize the need for more adequate tooling support in the legal domain. The individual phases in the workflow (see 5) need different types of support for these challenges. As discussed before (see 2), some of the challenges align with standard practices in text-based research domains, which allows us to reuse some existing concepts, but with the need to adapt and integrate them into the specific needs of law scholars.
@@
We identify three key areas of VA that hold significant benefits for jurisprudence: (1) data navigation: for effective exploration and cross-connection of document collections, (2) knowledge representation: for properly structuring data and metadata while integrating domain knowledge with explicit knowledge from legal entities, and (3) analytical reasoning: for synthesizing and documenting findings with the support of AI. In 2, we showcase the visual components of an exemplary traditional user interface of a commonly used German legal database. Alongside this traditional UI, we propose potential enhancements through VA techniques. The traditional keyword-based search (see 4) consists of three parts: (1) the formal legal structure of the inspected law for data navigation as a hierarchical list, (2) the display of the legal norm’s contents as a main body of text, and (3) a simple list of related legal documents. Based on the literature and our findings, the following enhancements are possible. On the left-hand side, in addition to the hierarchical list, the UI (1) visualizes the hierarchy of the Bundespolizeigesetz (Federal Police Act) using a collapsible tree structure as a conceptual tree map. It contains the law’s formal structure labeled by its section titles. Users can navigate the law through interaction with the structure, moving between legal norms. The UI (2) renders the selected legal norm’s contents at the center, representing explicit knowledge. The rendering highlights the keywords from the query in the contents verbatim. The user interface backlinks to their documents for explicitly referenced legal entities, simultaneously displaying a preview on hover. On the right-hand side, the legal database (3) lists legal documents related to the currently displayed document, like legal norms, law commentary, and court rulings. For the three key areas of VA discussed above, we analyze the deficiencies, explore the state-of-the-art literature, and distill possible improvements in jurisprudence through VA, while also discussing the challenges at this intersection.
@@ 6.1 Data Navigation
In VA, data navigation encompasses the exploration of and the movement in information spaces [shrinivasanSupportingAnalyticalReasoning2008]. For jurisprudence, this translates to navigating formally structured legal document collections. The traditional UI of legal databases (refer to 2) relies on a collapsible tree-based list to represent hierarchies. However, this approach fails to provide semantic context during navigation. Since legal structures can be nested several levels deep [bundesministeriumfuerjustizHandbuchRechtsfoermlichkeit2024, p. 152], their tree representation can become cumbersome. Additionally, tree-based lists inadequately support the simultaneous navigation of multiple legal structures – a common requirement in jurisprudential research.
@@
[merklExplorationLegalText1997] introduced the concept of navigating such corpora using a Hierarchical Feature Map [merklExplorationLegalText1997], a Treemap variant that organizes clusters across layers representing the hierarchy. Each layer corresponds to a hierarchical level, with subsequent layers revealing sub-clusters. Despite their improvement, the static, semi-automatically generated maps require manual input for cluster titles, limiting scalability. Building on this idea, [lettieriLegalMacroscopeExperimenting2017] started employing Treemaps [johnsonTreeMapsSpaceFillingApproach1991] in the Legal Doctrine Semantic Navigator for navigating formally structured legal documents through drill-down and roll-up operations [lettieriLegalMacroscopeExperimenting2017]. Later, they introduced the sliding Treemap, a variation, enabling the navigation of rulings by the European Court of Justice and other legal documents on mobile devices [lettieriAffordanceLawSliding2020]. Treemaps with interactive drill-downs can display multiple paths simultaneously across hierarchies and can visually encode guidance by enlarging documents of higher relevance [lettieriLegalMacroscopeExperimenting2017] (refer to 2, a).
@@
Beyond Treemaps, Icicle Plots [kruskalIciclePlotsBetter1983] have emerged as an effective visualization method for hierarchical data. DH researchers employ these plots for document comparison [tytarenkoHierarchicalTopicMaps2024]. [kimArchiTextInteractiveHierarchical2020] use Icicle Plots to visualize a hierarchy of topics during their interactive modeling [kimArchiTextInteractiveHierarchical2020], while [tytarenkoHierarchicalTopicMaps2024] also visualize a hierarchy of topics to compare documents [tytarenkoHierarchicalTopicMaps2024] (refer to 2, a).
@@
We propose a hybrid approach combining the strengths of both visualization techniques based on the following observations. Since Treemaps perform worse at displaying large hierarchies than Icicle Plots [macquistenEvaluationHierarchicalVisualization2020], distributing the visual representation between these complementary visualizations is beneficial. A Treemap would display the first few layers of the currently selected path as an overview, while a juxtaposed Icicle Plot would present the remainder of the path in the hierarchy. Employing the two visualizations in parallel mitigates scalability issues and leverages the benefits of interactively linking both visualizations [burchPowerInteractivelyLinked2022], offering complementary perspectives on legal corpora. This combination would support legal scholars during the initial workflow phase (refer to $W_{\texttt{P1}}$), facilitating the discovery and scoping of legal document collections.
@@
Alternative hierarchical visualization techniques, such as Bubble Treemaps or Sunbursts [macquistenEvaluationHierarchicalVisualization2020] can also be employed but perform worse for small hierarchies than our proposed approaches. Still, we emphasize the need for exploring other design alternatives such as circular grid layouts [qiuVADISVisualAnalytics2025] and hierarchical edge bundling [holtenHierarchicalEdgeBundles2006] for radial visualizations like a radial dendrogram. Though less prevalent in the legal domain, these alternatives are worth considering for expanding the design space.
@@ 6.2 Knowledge Representation
The interactions with visualizations for data navigation heavily depend on the user’s prior knowledge, explicit and tacit [wangDefiningApplyingKnowledge2009]. In jurisprudence, explicit knowledge can manifest as connections between legal documents [lacavaLawNetVizWebbasedSystem2022], where one norm references another legal norm, establishing a relationship. The traditional UI of legal databases displays these relationships through backlinks, with a pop-up and metadata, thereby representing explicit knowledge. The tree-based list of the law’s hierarchy (left) and the list of related legal documents (right) also express explicit knowledge contained in legal documents. Meanwhile, legal scholars’ tacit knowledge about the norm application is internalized as part of legal scholarship. However, it is not articulated [ruethersRechtstheorieUndJuristische2022], which would be especially important for legal reasoning (see 5.2).
@@
The literature proposes visualizations such as stacked graphs and geographic maps, which have been employed to visualize the metadata of legal documents [gomez2015understanding]. Similarly, [resckLegalVisExploringInferring2023] visualize legal structure by semantically segmenting it by paragraphs [resckLegalVisExploringInferring2023]. Meanwhile, other works statically analyze the complexity by visualizing explicit references between articles [bokwonleeNetworkStructureReveals2018]. Beyond jurisprudence, explicit knowledge of text-based documents is visualized in the context of discussion forum posts [jian-syuanwongMessageLensVisualAnalytics2018] and document comparison [tytarenkoHierarchicalTopicMaps2024]. However, these examples often neglect to facilitate the articulation of tacit knowledge. Instead, VA should consider tacit knowledge [federicoRoleExplicitKnowledge2017]. For instance, [wagnerKnowledgeassistedVisualMalware2017] allow cybersecurity experts to classify malware behavioral rules [wagnerKnowledgeassistedVisualMalware2017], while [mistelbauerSmartSuperViews2012] enhance medical visualizations by ranking views based on experience [mistelbauerSmartSuperViews2012]. [endertSemanticInteractionVisual2012] define several semantic interactions that support analytical reasoning for text analytics [endertSemanticInteractionVisual2012]. The authors also provide an implementation, letting users determine the importance of phrases through highlighting. Accordingly, documents containing the exact phrase become closer in a spatial layout (refer to 2, b). Similarly, a diagram of concepts connected by labeled arcs constitutes a Concept Map, where the arcs created by semantic interactions describe tacit knowledge [canasConceptMapsIntegrating2005] (refer to 2, b).
@@
Translating these solutions from VA to jurisprudence, we observe that visually laying out legal concepts in a map is an established approach [burkhardtVisualLegalAnalytics2019, ginevraperuginelliKnowledgeLawBig2019, onamiLegalVizLegalText2025]. Concept maps can enhance VA for jurisprudence by converting legal scholars’ domain knowledge into explicit knowledge through semantic interactions, thereby improving the effectiveness of LLMs. With KMTLabeler, [wangKMTLabelerInteractiveKnowledgeAssisted2024] propose a labeling tool for medical texts integrating VA with LLMs [wangKMTLabelerInteractiveKnowledgeAssisted2024], allowing experts to express their domain knowledge by keyword labeling rules. We can also connect the knowledge representation to legal reasoning since a Concept Map resembles a network. Both support the latter two phases of the jurisprudence workflow (refer to $W_{\texttt{P2}}$ and $W_{\texttt{P3}}$).
@@
While networks are suitable for supporting this workflow phase, node-link diagrams may not be the most effective visualization method. Matrices, i.e., tabular displays of graphs, are proven methods for visualizing network data, but are underutilized in VLA. Hence, studying hybrid approaches and alternative representations specific to the legal domain can enrich the VLA design space.